OpenStudy (frank_11):

What is the answer for |x-8|=|x+10| And how?

OpenStudy (trojanpoem):

either x-8 = x + 10 (same as -(x-8) = -(x+10)) 18 = 0 Refused solution. or x -8 = -(x+10) (consider that -(x-8) = x+10 is equivalent to this) x - 8 = -x - 10 2x = -2 x = -1

OpenStudy (frank_11):

Thank you